g(x)=4x^2-3x+1 slove it please
Answer by checkley77(12569)

g(x)=4x^2-3x+1
g(x)=(4x+1)(x-1)
4x+1=0
4x=-1
x=-1/4 ans.
x-1=0
x=1 ans.